OpenStudy (twizttiez):

Solve, using the substitution method. y – 2x = 8 16 + 4x = 2y A. The only solution is (1, 10). B. There is no solution. C. There are an infinite number of solutions. D. The only solution is (24, 0).

OpenStudy (twizttiez):

@Loser66

OpenStudy (anonymous):

Okay. You need to take one of those equations and change it into either y= or x= so you can plug it in to the OTHER equation.

OpenStudy (twizttiez):

Ok

OpenStudy (anonymous):

or, "substitute it".

OpenStudy (anonymous):

I would suggest turning y-2x=8 into y=2x+8

OpenStudy (anonymous):

Then, turning around and plugging it into the other equation, like so. 16+4x=2(2x+8)

OpenStudy (anonymous):

Distribute the 2 and solve.

OpenStudy (twizttiez):

Ok

OpenStudy (twizttiez):

I got A is that correct?

OpenStudy (twizttiez):

@mikaylabean

OpenStudy (anonymous):

Okay, not quite. Let's work through It together?

OpenStudy (twizttiez):

Ok

OpenStudy (anonymous):

So, I left you with 4x+8=16+4x

OpenStudy (anonymous):

First, subtract the 8 from both sides.

OpenStudy (twizttiez):

Ok

OpenStudy (anonymous):

What are you left with?

OpenStudy (twizttiez):

4x+4=16+0x?

OpenStudy (anonymous):

4x+4=16

OpenStudy (twizttiez):

:/ i knew that

OpenStudy (anonymous):

Not quite. 4x=8+4x

OpenStudy (twizttiez):

Ok so then what do i do?

OpenStudy (anonymous):

You're subtracting the 8 from the left side from both sides. Next, you need to subtract 4x from both sides, youll be left with 0=8 and thats not true, so there are no solutions

OpenStudy (twizttiez):

Ohhhhh ok could you help with a few more?

OpenStudy (anonymous):

Sure thing.

OpenStudy (twizttiez):

YAY! so kawai

OpenStudy (twizttiez):

One moment lemme open up a new ?

OpenStudy (anonymous):

Okay.